# Lite Any Stereo: Efficient Zero-Shot Stereo Matching

### v1

@@ 4.1 Benchmarks, Metrics, and Baselines
Benchmarks. We evaluate our method on five widely-used real-world stereo datasets. Middlebury [middlebury] is an indoor dataset containing 15 stereo pairs with high-quality ground truth captured using structured light. In the main paper, we report results under half-resolution and non-occluded settings; full-resolution and quarter-resolution results are provided in the supplementary materials. ETH3D [eth3d] consists of 27 grayscale stereo pairs with laser-scanned ground truth, covering both indoor and outdoor scenes. KITTI 2012 [kitti12] and KITTI 2015 [kitti15] include 194 and 200 stereo pairs, respectively, captured in outdoor driving environments, with ground truth obtained via LiDAR. DrivingStereo weather [yang2019drivingstereo] contains driving scene images under four different weather conditions, where each class of weather contains 500 frames. We report results at full resolution. Note that none of these datasets are used in the training process.
@@ 4.2 Implementation Details
Lite Any Stereo is implemented using PyTorch. The model is trained for 150K, 50K, and 100K steps in stage 1, stage 2, and stage 3 with a total batch size of 176 on NVIDIA A100 GPUs. We adopt the AdamW optimizer [adam] with a one-cycle learning rate schedule, where peak learning rate is set to $2\times 10^{-4}$. During training, input images are randomly cropped to $256\times 512$, and then finetuned on $320\times 736$. The maximum disparity value $\text{D}_{\text{max}}$, is set to 192, following the same configuration in previous works [guo2024lightstereochannelboostneed, xu2025banet].
@@ 4.3 Ablation Study
We investigate various design choices and training strategies for our model. Unless otherwise specified, all models are trained for 100K iterations without data augmentation on a 1.4M-image subset of synthetic datasets [wen2025foundationstereozeroshotstereomatching, li2022practical, cabon2020vkitti2, fallingthings, mayer2016large]. We report D1 errors on KITTI 2012 [kitti12] (K.12) and KITTI 2015 [kitti15] (K.15), Bad 1.0 error on ETH3D [eth3d] (E.), and Bad 2.0 error on Middlebury [middlebury] (M.), as in Tab. 6(g).
Cost Aggregation Design. In Tab. 6(g)(a), we first compare different aggregation choices (Fig. 5) and observe that the 3D–2D hybrid design is effective. We then evaluate alternative 2D layer choices in Tab. 6(g)(c), and find that ConvNeXt layers [liu2022convnet] provide the best trade-off between accuracy and efficiency. Building upon the ConvNeXt-based 2D layers, we further evaluate the influence of 3D convolution kernel sizes (Tab. 6(g)(b)). Unlike the findings in [wen2025foundationstereozeroshotstereomatching], where larger kernels improved performance, our results show that standard $(3,3,3)$ configuration achieves the highest accuracy in our baseline. Finally, as shown in Tab. 6(g)(d), increasing the proportion of 3D blocks leads to performance degradation, particularly on the Middlebury dataset under limited MACs. Therefore, we retain only a small 3D component to preserve disparity perception while maintaining efficiency.
@@
Table 5: Results on KITTI 2012 [kitti12] and KITTI 2015 [kitti15] leaderboard. The MACs are measured for an input size of $1242\times 375$. Our finetuned version ranks 1st on the leaderboards among all efficient methods at the time of submission.
@@
DispNetC [mayer2016large] 4.11 4.65 2.77 3.20 0.9 / 1.0 4.32 4.41 4.34 -
DeepPruner-Fast [duggal2019deeppruner] - - - - - 2.32 3.91 2.59 194
AANet+ [xu2020aanet] 1.55 2.04 1.20 1.58 0.4 / 0.5 1.65 3.96 2.03 -
DecNet [yao2021decomposition] - - - - - 2.07 3.87 2.37 -
BGNet+ [xu2021bilateral] 1.62 2.03 1.16 1.48 0.5 / 0.6 1.81 4.09 2.19 77
HITNet [tankovich2021hitnet] 1.41 1.89 1.14 1.53 0.4 / 0.5 1.74 3.20 1.98 47
CoEx [bangunharcana2021correlate] 1.55 1.93 1.15 1.42 0.5 / 0.5 1.79 3.82 2.13 54
MobileStereoNet-2D [shamsafar2022mobilestereonet] - - - - - 2.49 4.53 2.83 127
MobileStereoNet-3D [shamsafar2022mobilestereonet] - - - - - 2.75 3.87 2.10 564
Fast-ACVNet [xu2022acvnet] 1.68 2.13 1.23 1.56 0.5 / 0.6 1.82 3.93 2.17 72
Fast-ACVNet+ [xu2023accurate] 1.45 1.85 1.06 1.36 0.5 / 0.5 1.70 3.53 2.01 85
Lite-CREStereo++ [Jing_2023_ICCV] 1.43 1.82 1.12 1.44 0.5 / 0.5 1.79 3.53 2.08 93
LightStereo-M [guo2024lightstereochannelboostneed] 1.56 1.91 1.10 1.36 0.5 / 0.5 1.81 3.22 2.04 33
LightStereo-L [guo2024lightstereochannelboostneed] 1.55 1.87 1.10 1.33 0.5 / 0.5 1.78 2.64 1.93 84
BANet-2D [xu2025banet] 1.38 1.79 1.01 1.32 0.5 / 0.5 1.59 3.03 1.83 36
BANet-3D [xu2025banet] 1.27 1.72 0.95 1.27 0.5 / 0.5 1.52 3.02 1.77 78
@@ 4.4 Evaluation
Zero-Shot Generalization. Tab. 3 presents a quantitative comparison of zero-shot generalization performance across four public datasets. In the first block of the table, we restrict our model training to stage 2 on SceneFlow to ensure a fair comparison. Under this setting, our method achieves superior performance on most metrics while requiring the fewest MACs, highlighting the effectiveness of both model design and training strategy. We then evaluate under a more realistic setting as in [wen2025foundationstereozeroshotstereomatching], where methods are allowed to train on any available dataset except the target domains. In this setup, our approach consistently outperforms the other methods by a large margin. Our model achieves comparable or better results than non-prior-based accurate method [wang2024selective] with less than 1% of its MACs. Fig. 7 presents qualitative comparison on in-the-wild, high-resolution (22K) images. Our method generates smooth yet accurate disparity maps across different real-world scenes, whereas the other methods struggle to preserve geometric details.
@@
In-Domain Performance. Tab. 4 reports the results on the DrivingStereo weather subset. Interestingly, although our model is only trained via knowledge distillation on general datasets, it surpasses the teacher model (FoundationStereo [wen2025foundationstereozeroshotstereomatching]) by a significant margin. This highlights the strong effectiveness of our proposed approach. Although in-domain performance is not the primary focus of this work, we also evaluate our model on the KITTI online leaderboard (test set) trained on [kittidepth] without using its original annotations. As can be seen in Tab. 5, our model achieves the highest accuracy among all published efficient methods across almost all metrics at the time of submission.
Effectiveness of the Training Strategy. To verify the generality of our training strategy, we apply it to two representative efficient methods, LightStereo-M [guo2024lightstereochannelboostneed] and BANet-2D [xu2025banet]. In Tab. 6, our strategy consistently improves both methods across multiple benchmarks, which shows its broad applicability to different architectures.
@@ 4.5 Limitations
Lite-CREStereo++ [Jing_2023_ICCV] 63 38 55 40
LightStereo-M [guo2024lightstereochannelboostneed] 24 22 27 21
LightStereo-L [guo2024lightstereochannelboostneed] 34 30 35 27
BANet-2D [xu2025banet] 58 30 63 27
BANet-3D [xu2025banet] 40 25 44 22

### v2

@@ 4.1 Benchmarks, Metrics, and Baselines
Benchmarks. We evaluate our method on five widely-used real-world stereo datasets. Middlebury [46] is an indoor dataset containing 15 stereo pairs with high-quality ground truth captured using structured light. In the main paper, we report results under half-resolution and non-occluded settings; full-resolution and quarter-resolution results are provided in the supplementary materials. ETH3D [47] consists of 27 grayscale stereo pairs with laser-scanned ground truth, covering both indoor and outdoor scenes. KITTI 2012 [14] and KITTI 2015 [40] include 194 and 200 stereo pairs, respectively, captured in outdoor driving environments, with ground truth obtained via LiDAR. DrivingStereo weather [72] contains driving scene images under four different weather conditions, where each class of weather contains 500 frames. We report results at full resolution. Note that none of these datasets are used in the training process.
@@ 4.2 Implementation Details
Lite Any Stereo is implemented using PyTorch. The model is trained for 150K, 50K, and 100K steps in stage 1, stage 2, and stage 3 with a total batch size of 176 on NVIDIA A100 GPUs. We adopt the AdamW optimizer [31] with a one-cycle learning rate schedule, where peak learning rate is set to $2\times 10^{-4}$. During training, input images are randomly cropped to $256\times 512$, and then finetuned on $320\times 736$. The maximum disparity value $\text{D}_{\text{max}}$, is set to 192, following the same configuration in previous works [20, 66].
@@ 4.3 Ablation Study
We investigate various design choices and training strategies for our model. Unless otherwise specified, all models are trained for 100K iterations without data augmentation on a 1.4M-image subset of synthetic datasets [61, 32, 6, 53, 38]. We report D1 errors on KITTI 2012 [14] (K.12) and KITTI 2015 [40] (K.15), Bad 1.0 error on ETH3D [47] (E.), and Bad 2.0 error on Middlebury [46] (M.), as in Tab. 6(g).
Cost Aggregation Design. In Tab. 6(g)(a), we first compare different aggregation choices (Fig. 5) and observe that the 3D–2D hybrid design is effective. We then evaluate alternative 2D layer choices in Tab. 6(g)(c), and find that ConvNeXt layers [36] provide the best trade-off between accuracy and efficiency. Building upon the ConvNeXt-based 2D layers, we further evaluate the influence of 3D convolution kernel sizes (Tab. 6(g)(b)). Unlike the findings in [61], where larger kernels improved performance, our results show that standard $(3,3,3)$ configuration achieves the highest accuracy in our baseline. Finally, as shown in Tab. 6(g)(d), increasing the proportion of 3D blocks leads to performance degradation, particularly on the Middlebury dataset under limited MACs. Therefore, we retain only a small 3D component to preserve disparity perception while maintaining efficiency.
@@
Table 5: Results on KITTI 2012 [14] and KITTI 2015 [40] leaderboard. The MACs are measured for an input size of $1242\times 375$. Our finetuned version ranks 1st on the leaderboards among all efficient methods at the time of submission.
@@
DispNetC [38] 4.11 4.65 2.77 3.20 0.9 / 1.0 4.32 4.41 4.34 -
DeepPruner-Fast [13] - - - - - 2.32 3.91 2.59 194
AANet+ [70] 1.55 2.04 1.20 1.58 0.4 / 0.5 1.65 3.96 2.03 -
DecNet [75] - - - - - 2.07 3.87 2.37 -
BGNet+ [64] 1.62 2.03 1.16 1.48 0.5 / 0.6 1.81 4.09 2.19 77
HITNet [52] 1.41 1.89 1.14 1.53 0.4 / 0.5 1.74 3.20 1.98 47
CoEx [1] 1.55 1.93 1.15 1.42 0.5 / 0.5 1.79 3.82 2.13 54
MobileStereoNet-2D [48] - - - - - 2.49 4.53 2.83 127
MobileStereoNet-3D [48] - - - - - 2.75 3.87 2.10 564
Fast-ACVNet [65] 1.68 2.13 1.23 1.56 0.5 / 0.6 1.82 3.93 2.17 72
Fast-ACVNet+ [68] 1.45 1.85 1.06 1.36 0.5 / 0.5 1.70 3.53 2.01 85
Lite-CREStereo++ [25] 1.43 1.82 1.12 1.44 0.5 / 0.5 1.79 3.53 2.08 93
LightStereo-M [20] 1.56 1.91 1.10 1.36 0.5 / 0.5 1.81 3.22 2.04 33
LightStereo-L [20] 1.55 1.87 1.10 1.33 0.5 / 0.5 1.78 2.64 1.93 84
BANet-2D [66] 1.38 1.79 1.01 1.32 0.5 / 0.5 1.59 3.03 1.83 36
BANet-3D [66] 1.27 1.72 0.95 1.27 0.5 / 0.5 1.52 3.02 1.77 78
@@ 4.4 Evaluation
Zero-Shot Generalization. Tab. 3 presents a quantitative comparison of zero-shot generalization performance across four public datasets. In the first block of the table, we restrict our model training to stage 2 on SceneFlow to ensure a fair comparison. Under this setting, our method achieves superior performance on most metrics while requiring the fewest MACs, highlighting the effectiveness of both model design and training strategy. We then evaluate under a more realistic setting as in [61], where methods are allowed to train on any available dataset except the target domains. In this setup, our approach consistently outperforms the other methods by a large margin. Our model achieves comparable or better results than non-prior-based accurate method [58] with less than 1% of its MACs. Fig. 7 presents qualitative comparison on in-the-wild, high-resolution (22K) images. Our method generates smooth yet accurate disparity maps across different real-world scenes, whereas the other methods struggle to preserve geometric details.
@@
In-Domain Performance. Tab. 4 reports the results on the DrivingStereo weather subset. Interestingly, although our model is only trained via knowledge distillation on general datasets, it surpasses the teacher model (FoundationStereo [61]) by a significant margin. This highlights the strong effectiveness of our proposed approach. Although in-domain performance is not the primary focus of this work, we also evaluate our model on the KITTI online leaderboard (test set) trained on [54] without using its original annotations. As can be seen in Tab. 5, our model achieves the highest accuracy among all published efficient methods across almost all metrics at the time of submission.
Effectiveness of the Training Strategy. To verify the generality of our training strategy, we apply it to two representative efficient methods, LightStereo-M [20] and BANet-2D [66]. In Tab. 6, our strategy consistently improves both methods across multiple benchmarks, which shows its broad applicability to different architectures.
@@ 4.5 Limitations
Lite-CREStereo++ [25] 63 38 55 40
LightStereo-M [20] 24 22 27 21
LightStereo-L [20] 34 30 35 27
BANet-2D [66] 58 30 63 27
BANet-3D [66] 40 25 44 22